please show steps

Problem 9. (5 points) Cowen met L sta) dr – 3 sie)dt = 2. and 1” siedz – 5. Compute[ 3f(x) dx = 0
Answer #1

en that fix) dx = 3 and fex) dx = 7 = [tex) dx = I fox) dx - fone) dre = 3 - 7 = -4. given that given that { fonda = 5 (ten)